OpenStudy (anonymous):

Which of the following statements describes one part of completing the square for x^2 + 4x = 32? a. Take the square root of 36 and add 2. b. Take the square root of 32 and subtract 2. c. Take the square root of 36 and subtract 2.

OpenStudy (campbell_st):

the answers relate to the solution to the equation by using the complete the square method so you need to have completed the square to start the process of solving so to make the left hand side a perfect square what needs to be added...?

OpenStudy (anonymous):

four needs to be added

OpenStudy (campbell_st):

great so the equation becomes \[x^2 + 4x + 4 =32 + 4 ~~~~or~~~~~(x + 2)^2 = 36\] so you are being asked, if you have the new equation above, what steps d you need to take to solve for x

OpenStudy (anonymous):

I need to apply the square root property and solve for x so that makes the equation turn into \[x+2=\pm \sqrt{36}\]

OpenStudy (anonymous):

So x=4 x=-8

OpenStudy (anonymous):

So would the answer be a. Take the square root of 36 and add 2.? because at one point I subtract minus two (making it plus?) from the square root of 36 (6)

OpenStudy (campbell_st):

that's correct so the 2 steps are 1. square root both sides 2. subtract 2 so just match that to the answer choices

OpenStudy (campbell_st):

I think its option C you took the square root of 36 then subtracted 2 from both sides

OpenStudy (anonymous):

ohh ok I get it thanks for the help